If you deposit $500 in a bank account that earns 6 percent per year, how much total interest will you have earned after the third year? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1

PV = $500

t = 3 years

r = 6% p.a.

FV = PV*(1+r)^t = 500*1.06^3

FV = $595.51

Total interest earned = FV-PV = 595.51 - 500 = $95.51